Derive an expression for heat produced in an appliance of resistance R when current is I passed through it for time T.Find power dissipated across 4 ohm resistor.Is reading of A1 and A2 different?
1
Expert's answer
2018-09-06T16:42:08-0400
An expression for heat
Q=I^2·R·T
Power dissipated across 4 ohm resistor
P=I^2·R=4·I^2

Answer: Q=I^2·R·T; P=4·I^2